OpenStudy (darkbluechocobo):

Help with checking work with Classifying Conics by Using the Discriminant

OpenStudy (darkbluechocobo):

Identify the value for C in the following equation that would make the conic section a parabola: 2x^2 − 4xy + Cy^2 + 5y − 3 = 0

OpenStudy (darkbluechocobo):

-4^2-4(2)(c) 16-8c=0 c=8/16?

OpenStudy (darkbluechocobo):

@triciaal Can you help me figure out what I did wrong. My answer doesn't match the multiple choice :( C=3/5 C = 2 C = 5 C = -2 C = -1

OpenStudy (darkbluechocobo):

oh wait -8c=-16 c=-16/8

OpenStudy (darkbluechocobo):

c=-2

OpenStudy (triciaal):

negative divided by negative is positive

OpenStudy (darkbluechocobo):

would be positive 2 s; i messed up

OpenStudy (triciaal):